Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2018-01-10

Observers: Filomena Bufano, Marco Berton 
TAT: Annalisa De Cia, Morgan Fraser
Support team: Stefan Taubenberger, Christian Vogl, Andreas Floers

Calibrations:  biases.  Dome arcs+flats taken during daytime are useless (see below)   

Weather Conditions: Clear Sky, No clouds, seeing~ 1.1" (UT 1.00), improving down to 0.7"/0.8". 

UT                  Target               Type                   Comment
==============================================================================================
While observing the standard, as first target, we found that two grisms of EFOSC were not aligned. Needed manual intervention.
In order to avoid an excessive time loss, we switched directly to SOFI, leaving a note for the daytime TO.
01.30     SOFI image quality test
01.44     AT2017jan            FollowUp                JHK 
03.02     AT2017jfa             FollowUp                JHK ; seeing~1.0" 
04.18     AT2016jbu            FollowUp               JHK 
05.35     AT2017ens           FollowUp               deep K; seeing~0.7"
06:19     AT2017jfs             FollowUp                JHK
07:35     AT2017ens           FollowUp               deep K, seeing~0.8"

Comments